# Many Hands Make Light Work: An LLM-based Multi-Agent System for Detecting Malicious PyPI Packages

### v1

@@ 1 Introduction
A variety of approaches have been developed to automatically analyze security threats, including static rule-based scanners (Ruohonen et al., 2021a), signature-driven classifiers (Chakraborty et al., 2020a), clustering methods (Liang et al., 2023), and behavior sequence models (Zhang et al., 2025b). While these methods achieve encouraging performance, they often fail in the presence of obfuscation, indirect API usage, or logic-level concealment. The recent success of large language models (LLMs) has opened new opportunities for analyzing and reasoning about software systems. Their ability to interpret code semantics through natural language prompts makes them attractive for tasks such as code summarization, defect detection, and vulnerability identification (Chen et al., 2021; Ahmad et al., 2021). Despite this potential, applications in security-critical domains such as malicious code detection in open-source repositories remain limited. Moreover, many existing attempts treat LLMs as monolithic black boxes, where a single model instance is tasked with multiple stages of analysis without meaningful modularity or transparency (Wang et al., 2023; Ohm et al., 2020). Empirical evidence on Retrieval Augmented Generation (RAG) further indicates that retrieval alone does not reliably improve the detection of malicious intent in PyPI packages (Ibiyo et al., 2025). Altogether, this motivates the development of an approach that can reliably surface file-level signals and provide auditable package-level decisions.
@@ 4.1 Research Questions
  • MPHunter (Liang et al., 2024) is a baseline approach using clustering techniques to group PyPI packages to identify outliers.
@@ 5 Experimental Results
$\triangleright$ The prediction results for LAMPS and MPHunter on D1 are shown in Table 2, which confirms that MPHunter obtains encouraging performance as claimed by the authors (Liang et al., 2024), i.e., it gets 95.86% as Accuracy, 0.963 as Precision, 0.955 as Recall, and 0.957 F1-score. Still, LAMPS is better as it gains more accurate predictions with respect to all the metrics, i.e.,, Accuracy is 97.77%, and all Precision, Recall, and F1-score are equal to 0.977.222These scores are equal thanks to the balanced dataset (D1).
@@ 5.3 RQ3: Is the deployment of multi LLMs-based agents advantageous with respect to only one agent?
To isolate the contributions of input curation and package-level aggregation, we further evaluate two constrained single-agent regimes. In the first regime (SA-Concat), the agent processes all extracted files in a deterministic order until the model’s context window is reached, at which point the remaining files are truncated. This concatenation-based approach has been considered in prior work on code analysis with LLMs (Bui et al., 2023; Ahmad et al., 2021). In the second regime (SA-TopK), the agent processes only the top $K$ files identified by the extractor as most relevant to package functionality and security; we set K=3K{=}3 to ensure that typical packages remain within the context window. Similar top-$K$ retrieval heuristics have been explored in LLM-augmented code search and software engineering tasks (Li et al., 2022). Both regimes used a fixed prompt instructing the model to produce a binary decision with a brief rationale; the full template is reported in the online appendix. Together, these baselines highlight the limitations of a monolithic LLM in multi-file scenarios where input length and ordering can affect detection outcomes.
@@ 6 Related Work
Learning-based strategies have recently been explored as more adaptive alternatives. Liang et al. proposed MPHunter, a clustering-based method that groups similar packages to identify anomalous outliers (Liang et al., 2021). This approach scales better than rule-based heuristics but remains sensitive to surface-level similarities; adversaries can still disguise malicious logic using obfuscation or mimicry of benign structures. More recently, Zhang et al. demonstrated a behavior-sequence model capable of detecting malicious packages across ecosystems such as npm and PyPI using a single model trained on behavioral signals (Zhang et al., 2025a). While this expands generalization, such methods depend heavily on the availability of representative execution traces. At the ecosystem level, Alfadel et al. (Alfadel et al., 2023) and Nguyen et al. (Nguyen et al., 2022) have highlighted that malicious packages not only threaten end users but also exploit weaknesses in dependency resolution and API recommendation systems, reinforcing the need for holistic supply chain defenses.
@@ 6.2 Malware and Vulnerability Detection with Machine Learning
Before the advent of LLMs, a large body of research investigated the use of classical and deep learning techniques for detecting malware and security vulnerabilities in source code. These approaches typically represent code through syntactic or semantic abstractions such as abstract syntax trees (ASTs), control-flow graphs (CFGs), or data-flow graphs (DFGs), which are then encoded into feature vectors for machine learning models. Early work applied traditional classifiers such as support vector machines (SVMs) and random forests to handcrafted code features, demonstrating moderate success in vulnerability prediction but limited scalability to unseen projects (Chakraborty et al., 2020b; Zhou et al., 2019).
@@
Parallel research has applied machine learning to malware detection at the binary and package level. Traditional malware classifiers leverage opcode sequences, API call graphs, or byte-level n-grams combined with supervised classifiers (Raff et al., 2018; Liang et al., 2021). More recent work integrates deep neural networks for end-to-end malware detection, though these methods remain vulnerable to evasion through obfuscation, packing, or adversarial perturbations. Studies such as EC2 (Chakraborty et al., 2020b) highlight that ensemble techniques can improve detection rates across malware families but still face challenges in interpretability and cross-platform generalization.
@@ 6.3 Software Supply Chain Security in Package Ecosystems
The problem of malicious packages is part of the broader challenge of securing the software supply chain. Recent incidents such as the SolarWinds compromise and large-scale injections of malicious dependencies in ecosystems like npm and PyPI have underscored the systemic risks that arise when attackers exploit trust relationships in package distribution networks (Ohm et al., 2020; Zimmermann et al., 2019). Research in this space has addressed multiple layers of the supply chain, including dependency management, vulnerability propagation, and ecosystem governance.
Empirical studies have shown that dependency networks in ecosystems such as PyPI, npm, and Maven expose substantial transitive risk: a single compromised package can affect thousands of downstream projects (Zimmermann et al., 2019; De Cicco et al., 2022). Taxonomies of supply chain attacks (Ohm et al., 2020) highlight that injection through public repositories is only one of several attack vectors, alongside compromised build pipelines, stolen credentials, and malicious updates. Detection strategies, therefore, include dependency risk scoring, provenance verification, and anomaly detection across package metadata and version histories (Torres et al., 2021).
@@ 6.4 LLM-based MAS in Software Engineering
Recent work has begun to explore the application of multi-agent systems powered by large language models in software engineering tasks. These systems decompose complex activities into coordinated agents that collaborate via structured prompts, often outperforming monolithic pipelines. For example, CoRe (Wang et al., 2024) is a collaborative MAS for code reviewer recommendation, where agents leverage semantic representations of pull requests and reviewer expertise to improve reviewer assignment. Metagente (Nguyen et al., 2025) employs a multi-agent prompt optimization framework for summarizing GitHub README files, demonstrating improvements in text quality via ROUGE-based feedback. Other frameworks such as AutoAgents (Li et al., 2023b), CAMEL (Li et al., 2023a), AgentVerse (Chen et al., 2023), and CrewAI (Taulli and Deshmukh, 2025) illustrate the broader potential of role-specialized agents for cooperative reasoning, planning, and task execution.
Applications of MAS in software engineering extend beyond recommendation and summarization. DocAgent (Yang et al., 2025) applies a multi-agent architecture to automated code documentation generation, while Arcadinho et al. (Arcadinho et al., 2024) demonstrate the use of agent collaboration for test generation and evaluation of LLM-powered tools. Ronanki et al. (Ronanki et al., 2023) investigate conversational agents for requirements elicitation, showing that task decomposition improves coverage and reduces ambiguity. Beyond task-specific contributions, several surveys have recently mapped this emerging space. He et al. (He et al., 2025) review LLM-based MAS across the software lifecycle and articulate open challenges. Chen et al. (Chen et al., 2025) and Liu et al. (Liu et al., 2024) provide complementary surveys of multi-agent architectures and methodologies, highlighting challenges in coordination, evaluation, and communication protocols. At the systems level, Lee et al. (Lee et al., 2025) examine strategies for reliable decision-making in MAS, comparing aggregation schemes such as majority voting, feedback loops, and arbitration. Verdecchia (Verdecchia, 2025) proposes SALLMA, a reference architecture for LLM-based multi-agent systems, emphasizing modular orchestration and contextual memory management.

### v2

@@ 1 Introduction
A variety of approaches have been developed to automatically analyze security threats, including static rule-based scanners (Ruohonen et al., 2021a), signature-driven classifiers (Chakraborty et al., 2020), clustering methods (Liang et al., 2023), and behavior sequence models (Zhang et al., 2025). While these methods achieve encouraging performance, they often fail in the presence of obfuscation, indirect API usage, or logic-level concealment. The recent success of large language models (LLMs) has opened new opportunities for analyzing and reasoning about software systems. Their ability to interpret code semantics through natural language prompts makes them attractive for tasks such as code summarization, defect detection, and vulnerability identification (Chen et al., 2021; Ahmad et al., 2021). Despite this potential, applications in security-critical domains such as malicious code detection in open-source repositories remain limited. Moreover, many existing attempts treat LLMs as monolithic black boxes, where a single model instance is tasked with multiple stages of analysis without meaningful modularity or transparency (Wang et al., 2023; Ohm et al., 2020). Empirical evidence on Retrieval Augmented Generation (RAG) further indicates that retrieval alone does not reliably improve the detection of malicious intent in PyPI packages (Ibiyo et al., 2025). Altogether, this motivates the development of an approach that can reliably surface file-level signals and provide auditable package-level decisions.
@@ 4.1 Research Questions
  • MPHunter (Liang et al., 2023) is a baseline approach using clustering techniques to group PyPI packages to identify outliers.
@@ 5 Experimental Results
$\triangleright$ The prediction results for LAMPS and MPHunter on D1 are shown in Table 2, which confirms that MPHunter obtains encouraging performance as claimed by the authors (Liang et al., 2023), i.e., it gets 95.86% as Accuracy, 0.963 as Precision, 0.955 as Recall, and 0.957 F1-score. Still, LAMPS is better as it gains more accurate predictions with respect to all the metrics, i.e.,, Accuracy is 97.77%, and all Precision, Recall, and F1-score are equal to 0.977.222These scores are equal thanks to the balanced dataset (D1).
@@ 5.3 RQ3: Is the deployment of multi LLMs-based agents advantageous with respect to only one agent?
To isolate the contributions of input curation and package-level aggregation, we further evaluate two constrained single-agent regimes. In the first regime (SA-Concat), the agent processes all extracted files in a deterministic order until the model’s context window is reached, at which point the remaining files are truncated. This concatenation-based approach has been considered in prior work on code analysis with LLMs (Choi and Lee, 2023; Ahmad et al., 2021). In the second regime (SA-TopK), the agent processes only the top $K$ files identified by the extractor as most relevant to package functionality and security; we set K=3K{=}3 to ensure that typical packages remain within the context window. Similar top-$K$ retrieval heuristics have been explored in LLM-augmented code search and software engineering tasks (Chen et al., 2023a). Both regimes used a fixed prompt instructing the model to produce a binary decision with a brief rationale; the full template is reported in the online appendix. Together, these baselines highlight the limitations of a monolithic LLM in multi-file scenarios where input length and ordering can affect detection outcomes.
@@ 6 Related Work
Learning-based strategies have recently been explored as more adaptive alternatives. Liang et al. proposed MPHunter, a clustering-based method that groups similar packages to identify anomalous outliers (Liang et al., 2021). This approach scales better than rule-based heuristics but remains sensitive to surface-level similarities; adversaries can still disguise malicious logic using obfuscation or mimicry of benign structures. More recently, Zhang et al. demonstrated a behavior-sequence model capable of detecting malicious packages across ecosystems such as npm and PyPI using a single model trained on behavioral signals (Zhang et al., 2023). While this expands generalization, such methods depend heavily on the availability of representative execution traces. At the ecosystem level, Alfadel et al. (Alfadel et al., 2023) and Nguyen et al. (Nguyen et al., 2022) have highlighted that malicious packages not only threaten end users but also exploit weaknesses in dependency resolution and API recommendation systems, reinforcing the need for holistic supply chain defenses.
@@ 6.2 Malware and Vulnerability Detection with Machine Learning
Before the advent of LLMs, a large body of research investigated the use of classical and deep learning techniques for detecting malware and security vulnerabilities in source code. These approaches typically represent code through syntactic or semantic abstractions such as abstract syntax trees (ASTs), control-flow graphs (CFGs), or data-flow graphs (DFGs), which are then encoded into feature vectors for machine learning models. Early work applied traditional classifiers such as support vector machines (SVMs) and random forests to handcrafted code features, demonstrating moderate success in vulnerability prediction but limited scalability to unseen projects (Chakraborty et al., 2020; Zhou et al., 2019).
@@
Parallel research has applied machine learning to malware detection at the binary and package level. Traditional malware classifiers leverage opcode sequences, API call graphs, or byte-level n-grams combined with supervised classifiers (Raff et al., 2018; Liang et al., 2021). More recent work integrates deep neural networks for end-to-end malware detection, though these methods remain vulnerable to evasion through obfuscation, packing, or adversarial perturbations. Studies such as EC2 (Chakraborty et al., 2020) highlight that ensemble techniques can improve detection rates across malware families but still face challenges in interpretability and cross-platform generalization.
@@ 6.3 Software Supply Chain Security in Package Ecosystems
The problem of malicious packages is part of the broader challenge of securing the software supply chain. Recent incidents such as the SolarWinds compromise and large-scale injections of malicious dependencies in ecosystems like npm and PyPI have underscored the systemic risks that arise when attackers exploit trust relationships in package distribution networks (Ohm et al., 2020). Research in this space has addressed multiple layers of the supply chain, including dependency management, vulnerability propagation, and ecosystem governance.
Empirical studies have shown that dependency networks in ecosystems such as PyPI, npm, and Maven expose substantial transitive risk: a single compromised package can affect thousands of downstream projects. Taxonomies of supply chain attacks (Ohm et al., 2020) highlight that injection through public repositories is only one of several attack vectors, alongside compromised build pipelines, stolen credentials, and malicious updates. Detection strategies, therefore, include dependency risk scoring, provenance verification, and anomaly detection across package metadata and version histories (Zahan, 2023).
@@ 6.4 LLM-based MAS in Software Engineering
Recent work has begun to explore the application of multi-agent systems powered by large language models in software engineering tasks. These systems decompose complex activities into coordinated agents that collaborate via structured prompts, often outperforming monolithic pipelines. For example, CoRe (Wang et al., 2024) is a collaborative MAS for code reviewer recommendation, where agents leverage semantic representations of pull requests and reviewer expertise to improve reviewer assignment. Metagente (Nguyen et al., 2025) employs a multi-agent prompt optimization framework for summarizing GitHub README files, demonstrating improvements in text quality via ROUGE-based feedback. Other frameworks such as CAMEL (Li et al., 2023), AgentVerse (Chen et al., 2023b), and CrewAI (Taulli and Deshmukh, 2025) illustrate the broader potential of role-specialized agents for cooperative reasoning, planning, and task execution.
Applications of MAS in software engineering extend beyond recommendation and summarization. DocAgent (Yang et al., 2025) applies a multi-agent architecture to automated code documentation generation, while Arcadinho et al. (Arcadinho et al., 2024) demonstrate the use of agent collaboration for test generation and evaluation of LLM-powered tools. Ronanki et al. (Ronanki et al., 2023) investigate conversational agents for requirements elicitation, showing that task decomposition improves coverage and reduces ambiguity. Beyond task-specific contributions, several surveys have recently mapped this emerging space. He et al. (He et al., 2025) review LLM-based MAS across the software lifecycle and articulate open challenges. Chen et al. (Chen et al., 2024) and Liu et al. (Liu et al., 2024) provide complementary surveys of multi-agent architectures and methodologies, highlighting challenges in coordination, evaluation, and communication protocols. At the systems level, Lee et al. (Lee et al., 2025) examine strategies for reliable decision-making in MAS, comparing aggregation schemes such as majority voting, feedback loops, and arbitration. Becattini (Becattini et al., 2025) proposes SALLMA, a reference architecture for LLM-based multi-agent systems, emphasizing modular orchestration and contextual memory management.